# Break-for-Make: Modular Low-Rank Adaptations for Composable Content-Style Customization

### v1

@@ 6. Applications and Discussions
We demonstrate the effectiveness and versatility of our technique on various applications, including content-style customization of various textures and portraits.
@@
Our technique enables synthesizing high-quality content with a wide variety of user-controlled textures and materials, which can be used for customized product visualization, digital content creation, or material design. We present results for different textures (knit texture, burlap texture, denim texture, and fabric texture) in Fig. 17. The visualized results indicate that our method is capable of customizing generation for a variety of textures while maintaining content consistency with the reference images. With our approach, designers can easily showcase their products with custom material and textile options tailored to customer preferences. Compared to traditional rendering pipelines requiring extensive modeling and material setup, our data-driven approach significantly streamlines this process.
@@
Our proposed method demonstrates considerable promise in addressing customized content and style fusion. However, it is imperative to acknowledge instances where the model sometimes exhibits style reference overfitting tendencies, leading to suboptimal results in certain scenarios. As presented in Fig. 19, the lack of diversity between style reference images leads to the observed overfitting. To address the challenge of overfitting, future iterations of our method will incorporate rigorous regularization techniques and enhanced data preprocessing methodologies. Additionally, the integration of cross-validation procedures and model simplification strategies will be explored to promote improved generalization performance.
Figure 19. Bad cases of our method. Sometimes the results tend to overfit into the style reference image.
@@
While our method performs well on content-style customization, generating images with complex or rare content/style only by using textual prompts remains challenging. Specifically, our method leverages the class priors in the T2I model when learning the content or style of given images ($e.g.$ “a [c1] dog” leverages “dog” as a class prior, “a [s1] yarn style” leverages “yarn” as a class prior) (Ruiz et al., 2023). When customized content or style images are highly complex or rare, obtaining accurate priors through simple prompts becomes challenging, leading to a decrease in the fidelity of generated images.

### v2

@@ 6. Applications and Discussions
We demonstrate the effectiveness and versatility of our technique across various applications, including content-style customization of diverse textures and portraits.
@@
Our technique enables the synthesis of high-quality content with a wide variety of user-controlled textures and materials, which can be leveraged for customized product visualization, digital content creation, or material design applications. We present results for different textures (knit texture, burlap texture, denim texture, and fabric texture) in Fig. 17. The visualized results indicate that our method is capable of customizing generation for a diverse range of textures while maintaining content consistency with the reference images. With our approach, designers can easily showcase their products with custom material and textile options tailored to customer preferences. Compared to traditional rendering pipelines requiring extensive modeling and material setup, our data-driven approach significantly streamlines this process.
@@
While our proposed method demonstrates considerable promise in addressing customized content and style fusion, it is essential to acknowledge instances where the model occasionally exhibits an influence from the style reference on the background regions of the generated images, as presented in Fig. 19. This observed influence on the background regions in the generated results can be attributed to the limited diversity among the style reference images. To mitigate this issue, future iterations of our method will incorporate rigorous regularization techniques and enhanced data preprocessing methodologies. Furthermore, the integration of cross-validation procedures and model simplification strategies will be explored to promote improved generalization performance.
Figure 19. Bad cases of our method. Sometimes style reference can induce undesirable influences on the background generation in the output results.
@@
While our method performs well on content-style customization, generating images with complex or rare content/style solely by using textual prompts remains a challenging task. Specifically, our method leverages the class priors in the T2I model when learning the content or style of given images ($e.g.,$ “a [c1] dog” leverages “dog” as a class prior, “a [s1] yarn style” leverages “yarn” as a class prior) (Ruiz et al., 2023). When the customized content or style images are highly complex or rare, obtaining accurate priors through simple prompts becomes challenging, leading to a decrease in the fidelity of the generated images.